An audio speaker producing a steady sound at an outdoor concert is 18 ft away from you. If you move to a position where the speaker is 67 ft distant, by what factor will the amplitude of the sound change?



1
Expert's answer
2022-04-18T08:50:19-0400

The  amplitude of the sound is inversely proportional to the distance

"A\sim\frac{1}{r}"

Hence

"\frac{A_2}{A_1}=\frac{r_1}{r_2}=\frac{18}{67}=0.27"

Answer: 0.27